# Topologically protected synchronization in networks

### v1

@@ Abstract
In a graph, we say that two nodes are topologically equivalent if their sets of first neighbors, excluding the two nodes, coincide. We prove that nonlinearly coupled oscillators located on a group of topologically equivalent nodes can get easily synchronized when the group forms a fully connected subgraph (or combinations of these), regardless of the status of all the other oscillators. More generally, any change occurring in the inner part of the remainder of the graph will not alter the synchronization status of the group. Typically, the group can synchronize when $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$, $k^{(\mathrm{IN})}$ and $k^{(\mathrm{OUT})}$ being the common internal and outgoing degree of each node in the group, respectively. Simulations confirm our analysis and suggest that groups of topologically equivalent nodes play the role of independent pacemakers.
@@
Many physical, biological, chemical, and technological systems can effectively be seen as networks of interacting oscillators, each with its own natural frequency. Two oscillators are considered neighbors if they are coupled by some nonlinear function of their phase difference. Spontaneous synchronization represents one of the most intriguing and ubiquitous aspects of these systems StrogatzC ; Pikovsky ; ArenasReview and has countless applications such as, pacemaker cells in the heart Heart , pacemaker cells in the brain Brain , flashing fireflies Fireflies , arrays of lasers Lasers , and superconducting Josephson junctions Josephson . Put simply, if the coupling constant $J$ is larger than the spread $\Delta\omega$ of the natural frequency distribution in the system, a finite portion of its oscillators tends to synchronize, i.e., they tend to rotate according to a common mean frequency and their phases fall into step with one another (locked in phase). Kuramoto-like models, where the coupling between two oscillators is modulated by the sine of their phase difference, provide the most popular approach for addressing synchronization in complex systems. In its mean-field version, where each oscillator is coupled to all other oscillators, the model allows to be exactly solved Strogatz2000 while, in the general case, despite no exact solution is known, rigorous bounds Strogatz1988 , large scale simulations and theoretical studies ArenasReview ; Munoz ; Kahng have succeeded to provide a general description of the different synchronization scenarios that can take place as a function of the topology of the underlying graph Goltsev , ranging from regular to random, small-world, and even scale-free. For the present manuscript, we emphasize the role played by modularity. Many complex networks are modular, i.e., can be seen as subgraphs (also known as groups or communities) with different internal and external connectivities Barabasi_Hierarchical ; Santo . It has been understood that, in modular networks, the densely connected communities synchronize first, and subsequently, the larger and less densely connected ones also tend to synchronize until full synchronization is achieved ArenasPRL ; ArenasReview .
Most of the above literature analyzes synchronization in the thermodynamic limit, viewed as a phase transition characterized by an emergent phase diagram containing essentially an incoherent region (where most of the oscillators follow their own natural frequencies) and a coherent one (where most of the oscillators get locked within each other). Here, we consider a rather different issue related to a local and protected synchronization: “local” because it concerns a finite group of oscillators embedded in some arbitrary graph containing the group; “protected” because the remainder of the graph cannot affect the synchronization status of the group, even when the former is subject to noise. Formally, our work belongs to what is known as synchronization of chaotic systems via bidirectional coupling Boccaletti , a branch less developed than the unidirectional case (or “slave-master” Note Boccaletti2 ; Topological_Non_Hermitian . At any rate, to the best of our knowledge, our analysis and findings — based on topological equivalence combined with modularity — have never been reported. In recent years, topology has been shown to play a major role in many areas of physics, ranging from topological insulators Topological_Insulator to quantum computation Topological_Quantum_Computing ; in this sense, synchronization cannot be an exception. We stress however that, despite a reminiscent terminology, our definition of topologically equivalent (TE) nodes is quite unrelated to these works.
@@
Given a graph with $N$ nodes and symmetric adjacency matrix $\bm{a}$, consider a Kuramoto-like model on the top it, where the phases of $N$ oscillators, $\theta_{1},\ldots,\theta_{N}$, with natural frequencies $\omega_{1},\ldots,\omega_{N}$, evolve according to
$\dot{\theta}_{i}=\omega_{i}+J\sum_{j}a_{i,j}\sin\left(\theta_{j}-\theta_{i}\right),$ (1)
$J>0$ being a coupling constant.
@@
where we have introduced the two functions
@@
Let us now suppose that node 1 and node 2 “see” the same remaining graph, i.e., i.e.,
@@
We shall say that the two nodes are TE. In this case
@@
where ${\theta}^{*}\in(\min\{\theta_{1},\theta_{2}\},\max\{\theta_{1},\theta_{2}\})$ and we have applied the mean-value theorem to the function $h(\theta)$. Note that ${\theta}^{*}$ is itself an unknown function of $\theta_{1}$ and ${\Theta}$ (or, alternatively, $\theta_{2}$ and ${\Theta}$), yet, as we shall see in a moment, it is worth to consider Eq. (7). We will use the following bounds:
$h^{\prime}(\theta)\leq k^{(\mathrm{OUT})}\quad\forall\theta,$ (8)
@@
$-\theta\leq-\sin(\theta)<-\theta(1-\gamma),\quad 0\leq\theta<1,$ (10)
where
@@
Equal frequencies with $a_{1,2}=1$. Let us assume that the two nodes have also equal natural frequencies, $\omega_{1}=\omega_{2}$. From Eq. (7) we see that, in the manifold $\theta_{1}=\theta_{2}$ ($=\theta^{*}$), the two external fields cancel out, in other words, ${\Theta}=~{}0$ provides a fixed point of the equation. Let us introduce $\lambda=k^{(\mathrm{OUT})}-2(1-\gamma)$ and $\mu=k^{(\mathrm{OUT})}+2$. From Eq. (7), by using the bounds (8) and (10), we see that, if $1>{\Theta}\geq 0$,
$-J\mu{\Theta}\leq\dot{{\Theta}}\leq J\lambda{\Theta},$ (12)
@@
while, if $-1<{\Theta}<0$, hold the opposite inequalities. Let us assume that, for the initial condition $\Theta_{0}=\Theta(t=0)$, we have $10<\Theta_{0}<$ 1. By continuity, there exists a sufficiently small time $t_{1}$ such that $01>{\Theta}(t)>$ 0 for $t\in[0,t_{1})$ whereby, from Eq. (12),
${\Theta}_{0}e^{-J\mu t}\leq{\Theta}\leq{\Theta}_{0}e^{J\lambda t}.$ (13)
@@
On the other hand, if $\lambda<0$, Eq. (13) and the fact that each $\theta_{i}(t)$ (an hence also $\theta_{2}-\theta_{1}$) as well as $\dot{\theta}_{i}(t)$ are continuous, imply that $t_{1}=\infty$, so that Eq. (13) holds for any $t$ SM . For simplicity, in the subsequent cases, on assuming certain constrains on the initial conditions, we shall limit ourselves to check that the upper bounding solution satisfies the same constrains for any $t$. Note however that, in general, simply bounding $|\Theta|$ is not enough for claiming good synchronization; the conservation of the sign of $\Theta$ is crucial.
@@
In conclusion, within a basin of attraction for the initial conditions contained in the region $|{\Theta}_{0}|<1$, a sufficient condition for the fixed point ${\Theta}=0$ to be stable is $k^{(\mathrm{OUT})}<2(1-\gamma)$. Of course, since $21<2(1-\gamma)<$ 2 and $k^{(\mathrm{OUT})}$ is integer, it follows that $k^{(\mathrm{OUT})}$ can be either 0 or 1, but, in view of generalizations to subsystems with $N^{\prime}>2$ TE oscillators, it is useful to keep in mind the inequality. Note that this synchronization between node 1 and node 2 occurs regardless of the dynamics of all the other oscillators, which in particular do not need to be synchronized. Note also that we have imposed only the equality of the frequencies of the two nodes, but the other nodes can have arbitrary frequencies and the result does not change. We stress that this is a consequence of the topological equivalence (4); when (4) does not hold, $h_{1}$ and $h_{2}$ remain two different functions so that, in the manifold $\theta_{1}=\theta_{2}$, they do not cancel out in Eq. (2). We can even imaging to modify the inner part of the remainder of the graph by dynamically removing, adding, or rewiring some of its links, as well as by allowing for the presence of any site-dependent noise: as far as such links are not those arriving at nodes 1 and 2, and as far as such noise applies only to the other nodes, the sufficient condition $k^{(\mathrm{OUT})}\leq 2(1-\gamma)$ remains satisfied, i.e., the subsystem fixed point $\theta_{1}=\theta_{2}$ keeps being stable.
@@
Equal frequencies with $a_{1,2}=0$. Before analyzing the most general case, it is worth also considering the sub-case in which the two nodes are disconnected,i.e., $a_{1,2}=0$ and have equal frequencies. In this case, the evolution equation for ${\Theta}$ is of no help because $\theta_{1}=\theta_{2}$ is no longer a stable attractor, in other words, in general, they do not get synchronized. However, from Eq. (1) we see that, if nodes 1 and 2 are TE, $\theta_{1}$ and $\theta_{2}$ obey the same identical equation: we simply have that one oscillator follows the other along the same identical trajectory and, in particular, if they started with the same initial condition, they will remain identical at any instant. In this specific case, ${\Theta}=0$ is no longer an attractor but rather a constant of the motion.
@@
Different frequencies with $a_{1,2}=1$. Let us now allow for the two frequencies to be different. If the pair is isolated, i.e., $k^{(\mathrm{OUT})}=0$, we have $h^{\prime}(\theta^{*})=0$ and Eq. (7) returns the known critical condition for the synchronization of $N^{\prime}=2$ oscillators: $2J>|\omega_{2}-\omega_{1}|$, the stability condition of the fixed point ${\Theta}$ being $\cos({\Theta})>0$. In the general case, $h^{\prime}(\theta^{*})\neq 0$, we must observe that, although Eq. (7) admits (under suitable conditions) a formal fixed point, due to the fact that $\theta^{*}$ depends on $\theta_{1}$ (besides ${\Theta}$), such a fixed point is conditioned on the value of $\theta_{1}$, which in turn is a function of time, rendering this formal fixed point useless. Our approach here is different: we assume that both $|\omega|=|\omega_{2}-\omega_{1}|/J$ and the initial condition $|{\Theta}_{0}|=|\theta_{2}(t=0)-\theta_{1}(t=0)|$ are sufficiently small and look for a bounding solution by exploiting again the bounds (8) and (10). Let us suppose ${\Omega}=\omega_{2}-\omega_{1}>0$ and $10<{\Theta}_{0}<$ 1. By continuity, at small enough times, we also have $10<{\Theta}<$ 1 and from Eq. (7) we get
@@
Equation (14) implies that, if $\lambda=k^{(\mathrm{OUT})}-2(1-\gamma)<0$, then ${\Theta}$, with initial condition $10<{\Theta}_{0}<$ 1, will remain bounded as follows
@@
Note that, for $0<\Omega/J|\lambda|<\Theta_{0}$, all the above procedure turns out to be consistent with the required bound $10<{\Theta}<$ 1 for any $t$.
In conclusion, we have proven that, under the three conditions $k^{(\mathrm{OUT})}<2(1-\gamma)$ (equivalent to $k^{(\mathrm{OUT})}\leq 1$), $10<{\Theta}_{0}<$ 1, and $0<\Omega/J|\lambda|<\Theta_{0}$, we have $0\leq\Theta\leq\Omega/J|\lambda|$ when $t\to\infty$. In other words, the phases of the two oscillators get asymptotically close to each other and the larger is $J$ the closer they stay. Moreover, as in the case of equal frequencies, we see that this bound holds regardless of the status of all the other oscillators located on the remainder of the graph and, again, any change in it, cannot affect the synchronization of the pair.
@@
Generalization to $N^{\prime}>2$ TE oscillators. Let us now consider a subsystem of $N^{\prime}=3$ TE nodes, say nodes 1, 2 and 3. Besides being TE with respect to the remainder of the graph, we want them to be TE among each other; in particular, each of them must have the same number of links pointing to the other two TE nodes. For $N^{\prime}=3$ there exists only one possibility, the one where $k^{(\mathrm{IN})}=2$, i.e., the three nodes form a triangle. Here we have introduced $k^{(\mathrm{IN})}$ as the internal (common) connectivity of the group, i.e., $k^{(\mathrm{IN})}$ is the number of links emanating from a node of the group and pointing to other nodes of the same group. Note that, formally, both the subsystem $N^{\prime}=3$ and the one already seen case $N^{\prime}=2$ (where it was $k^{(\mathrm{IN})}=1$), are fully connected (FC) graphs with $N^{\prime}$ nodes. However, as Fig. 1 shows, when $N^{\prime}>3$, there exist more configurations in which the $N^{\prime}$ TE nodes can be arranged. In fact, the number of possible arrangements tends to grow exponentially with $N^{\prime}$, but with a smaller rate for $N^{\prime}$ odd (see also SM ).
Figure 1: Graph associated to model (1) with $N=22$ nodes and $L=44$ links (15 red-ticker and 39 black-thin). The graph contains five groups of TE nodes: one FC with $N^{\prime}=3$ (nodes 1, 2 and 3), two FC with $N^{\prime}=2$ (nodes 7 and 6, and nodes 15 and 16, respectively); one with $N^{\prime}=4$ forming a square (nodes 9, 10, 11, and 12); and one FC with $N^{\prime}=4$ (nodes 19, 20, 21, 22). In each group, the internal links are drawn as ticker (red). In the four FC groups we have $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$ while in the other group we have $k^{(\mathrm{OUT})}<k^{(\mathrm{IN})}$, hence, in each group, the TC for synchronization are satisfied.
We introduce some notation and point out a few crucial points. Given $N^{\prime}$ TE nodes with indices $1,2,\ldots,N^{\prime}$, we indicate their corresponding phase differences by ${\Theta}_{i,j}=\theta_{i}-\theta_{j}$ and frequency differences by ${\Omega}_{i,j}=\omega_{i}-\omega_{j}$. Observe that ${\Theta}_{j,i}=-{\Theta}_{i,j}$ and that the variables $\{{\Theta}_{i,j}\}$ with $j>i$, are not all independent. For example, for $N^{\prime}=3$ we have the constrain ${\Theta}_{3,1}={\Theta}_{3,2}+{\Theta}_{3,1}$. In general, given $N^{\prime}$, we can always write a system of $N^{\prime}-1$ independent equations involving $N^{\prime}-1$ independent variables. We shall also assume $0<{\Theta}_{i,j}(0)<1/(N^{\prime}-1)$ so that, at small enough times, we also have $0<{\Theta}_{i,j}(t)<1/(N^{\prime}-1)$. Our general strategy is to use the bounds (8), (10); if the found bounding solution keeps satisfying the above constrains for any time $t$, the procedure is consistent. The resulting bounding system can be written vectorially:
@@
where $\bm{{\Theta}}=({\Theta}_{2,1},{\Theta}_{3,2},\ldots,{\Theta}_{N^{\prime},N^{\prime}-1})^{T}$, $\bm{{\Omega}}=({\Omega}_{2,1},{\Omega}_{3,2},\ldots,{\Omega}_{N^{\prime},N^{\prime}-1})^{T}$, and $\bm{B}$ is a $(N^{\prime}-1)\times(N^{\prime}-1)$ matrix that depends on the parameters $N^{\prime}$, $k^{(\mathrm{OUT})}$, $k^{(\mathrm{IN})}$, and $\gamma$. Let $\lambda_{i}$ and $\bm{u}_{i}$ be the eigenvalues and normalized eigenvectors of $\bm{B}$, respectively. The coefficients $c_{i}(t)=\bm{{\Theta}}\cdot\bm{u}_{i}$ satisfy the decoupled system $dc_{i}(t)/dt\leq\bm{{\Omega}}\cdot\bm{u}_{i}+J\lambda_{i}c_{i}(t)$. As a consequence, each component of the vector $\bm{{\Theta}}$ remains bounded if all the eigenvalues of $\bm{B}$ are negative, leading to the TC (note that the eigenvalues do not depend on $J$ or $\bm{{\Omega}}$). For example, in the FC case $N^{\prime}=3$ we have

### v2

@@ Abstract
In a graph, we say that two nodes are topologically equivalent if their sets of first neighbors, excluding the two nodes, coincide. We prove that nonlinearly coupled heterogeneous oscillators located on a group of topologically equivalent nodes can get easily synchronized when the group forms a fully connected subgraph (or combinations thereof), regardless of the status of all the other oscillators. More generally, any change occurring in the remainder of the graph will not alter the synchronization status of the group. Typically, the group can synchronize when $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$, $k^{(\mathrm{IN})}$ and $k^{(\mathrm{OUT})}$ being the common internal and outgoing degree of each node in the group, respectively. Simulations confirm our rigorous analysis and suggest that groups of topologically equivalent nodes act as independent pacemakers.
@@
Many physical, biological, chemical, and technological systems can effectively be seen as networks of interacting oscillators, each with its own natural frequency. Spontaneous synchronization represents one of the most intriguing and ubiquitous aspects of these systems StrogatzC ; Pikovsky ; ArenasReview and has countless applications such as, pacemaker cells in the heart Heart , pacemaker cells in the brain Brain ; Brain2 , flashing fireflies Fireflies , arrays of lasers Lasers , and superconducting Josephson junctions Josephson . Put simply, if the coupling constant $J$ is larger than the spread $\Delta\omega$ of the system’s natural frequencies, a finite portion of its oscillators synchronizes, i.e., they rotate according to a common frequency and their phases fall into step with one another. The Kuramoto model (KM) Kuramoto1 ; Strogatz2000 it is to the theory of synchronization as the Ising model is to the theory of phase transitions. In its mean-field version, the KM allows to be exactly solved Strogatz2000 while, in general, despite no exact solution is known, rigorous bounds Strogatz1988 , simulations, and theoretical studies ArenasReview ; Munoz ; Kahng have succeeded to provide a general description of the different synchronization scenarios taking place as a function of the topology of the underlying graph Goltsev , ranging from regular to random, small-world, and scale-free. For the present manuscript, we emphasize the role played by modularity. Many complex networks are modular, i.e., can be partitioned into subgraphs, also known as communities, with different internal and external connectivities Barabasi_Hierarchical ; Santo . It has been understood that, if $J>\Delta\omega$, the densely connected communities synchronize first and, subsequently, the larger and less densely connected ones also tend to synchronize, until full synchronization is achieved ArenasPRL ; ArenasReview .
The above literature analyzes synchronization in the thermodynamic limit as a phase transition, with a phase where each oscillator tends to rotate with its own natural frequency, and a phase where most of the oscillators rotate with a common frequency.
Here, we consider a rather different issue related to a local and protected synchronization: “local” because it concerns a finite group of oscillators embedded in some arbitrary graph containing the group; “protected” because the remainder of the graph cannot affect the synchronization status of the group, even under noise. Given a graph with $N$ nodes, consider a group of $N^{\prime}\leq N$ nodes and split the set of neighbors of each of these $N^{\prime}$ nodes into two sets: those contained within the group and those external to the group (external sets). We say that the $N^{\prime}$ nodes are topologically equivalent (TE) if they have the same number of links and identical external sets. In this work, we prove that, in the KM, a group of TE nodes can synchronize when the group forms a fully connected (FC) subgraph, regardless of the status of the other oscillators. We provide the exact conditions under which this occurs. Typically, the group can synchronize when $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$, where $k^{(\mathrm{IN})}$ and $k^{(\mathrm{OUT})}$ are the common internal and outgoing degrees of each node in the group.
Formally, this work falls within synchronization of chaotic systems via bidirectional couplings Boccaletti , a less-developed branch compared to the unidirectional case, where a common driver acting on uncoupled oscillators can induce synchronization Teramae ; Feng ; Hata ; Boccaletti2 . In the bidirectional case, analytical studies mainly concern mean-field modelsStrogatz2000 ; Rosenblum1 ; Rosenblum2 ; however, Pecora et al. Pecora showed that clusters of identical coupled oscillators can synchronize when they exhibit specific graph symmetries. On the other hand, MacArthur et al. MacArthur demonstrated that such symmetries — contrary to simple models of graph theory, like the random graph or the Barabási-Albert model — are very common in real-world networks. Moreover, MacArthur showed that most of these symmetric clusters are symmetric cliques, matching our definition of FC TE nodes. More recently, the transition toward complete synchronization through intermediate synchronized clusters have received further attention and the general mechanism has become better understood, especially in models involving identical oscillators Lin ; Lambiotte ; Fiore ; Gambuzza ; Multilayer ; Boccaletti3 .
The novelty of our work lies in the derivation of rigorous bounds for the phase differences of a group (or cluster) of coupled oscillators governed by a generic Kuramoto model with heterogeneous frequencies. These bounds depend solely on the model parameters and are independent of the dynamical states of the other oscillators. Unlike the works cited above, where the identity of the oscillators is a fundamental assumption for the analytical treatment, our KM-based analysis allows heterogeneous frequencies and supports multiple coexisting synchronized manifolds (more precisely, nearly-synchronized manifolds Note2 ), as observed in more realistic networks, particularly in biological systems.
@@
We derive our result starting from the simplest group and then generalize. Given a graph with $N$ nodes and symmetric adjacency matrix $a_{i,j}$, consider a KM on the top of it with coupling $J>0$: the phases of $N$ oscillators, $\theta_{1},\ldots,\theta_{N}$, with natural frequencies $\omega_{1},\ldots,\omega_{N}$, evolve according to
$\dot{\theta}_{i}=\omega_{i}+J\sum_{j}a_{i,j}\sin\left(\theta_{j}-\theta_{i}\right).$ (1)
@@
where we have introduced the two functionsNote3
@@
Let us now suppose that node 1 and node 2 are TE, i.e., they “see” the same remaining graph:
@@
In this case
@@
where ${\theta}^{*}\in(\min\{\theta_{1},\theta_{2}\},\max\{\theta_{1},\theta_{2}\})$ and we have applied the mean-value theorem to the function $h(\theta)$. Note that $\theta$$^{*}$ is itself an unknown function of $\theta_{1}$ and $\Theta$} (or, alternatively, $\theta_{2}$ and $\Theta$}), yet, as we shall see in a moment, it is worth to consider Eq. (7). We will use the bounds:
${-k^{(\mathrm{OUT})}\leq}h^{\prime}(\theta)\leq k^{(\mathrm{OUT})},\quad\forall\theta,$ (8)
@@
$-\theta<-\sin(\theta)<-\theta(1-\gamma),\quad 0<\theta<1,$ (10)
where SM
@@
Equal frequencies with $a_{1,2}=1$. Let us assume that the two nodes have equal frequencies, $\omega_{1}=\omega_{2}$. From Eq. (7) we see that the manifold $\theta_{1}=\theta_{2}$ ($\theta^{*}$), i.e., ${\Theta}=~0$, provides a fixed point. Let us introduce $\lambda=k^{(\mathrm{OUT})}-2(1-\gamma)$ and $\mu=k^{(\mathrm{OUT})}+2$. From Eq. (7), by using the bounds (8) and (10), we see that, if $1>{\Theta}>0$,
$-J\mu{\Theta}<\dot{{\Theta}}<J\lambda{\Theta},$ (12)
@@
while, if $1<{\Theta}<0$, hold the opposite inequalities. Let us assume that, for the initial condition $\Theta_{0}=\Theta(t=0)$, we have $0<\Theta_{0}<1$. By continuity, there exists a sufficiently small time $t_{1}$ such that $0<{\Theta}(t)<1$ for $t\in[0,t_{1})$ whereby, from Eq. (12), for $t\in[0,t_{1})$ we get
${\Theta}_{0}e^{-J\mu t}<{\Theta}<{\Theta}_{0}e^{J\lambda t}.$ (13)
@@
On the other hand, if $\lambda<0$, Eq. (13), and the fact that $\theta_{i}(t)$’s and $\dot{\theta}_{i}(t)$’s are continuous, imply that $t_{1}=\infty$. One can check this by a graphical construction, but also analytically SM . For simplicity, in the subsequent cases, on assuming certain constrains on the initial conditions, we shall limit ourselves to check that the upper bounding solution satisfies the same constrains for any $t$. Note however that, in general, simply bounding $|\Theta|$ is not enough for claiming good synchronization; the conservation of the sign of $\Theta$ is crucial.
@@
In conclusion, within a basin of attraction for the initial conditions $|{\Theta}_{0}|<1$, a sufficient condition for the fixed point ${\Theta}=0$ to be stable is $k^{(\mathrm{OUT})}<2(1-\gamma)$. Of course, since $1<2(1-\gamma)<2$ and $k^{(\mathrm{OUT})}$ is integer, it follows that $k^{(\mathrm{OUT})}$ can be either 0 or 1, but, in view of generalizations to subsystems with $N^{\prime}>2$ TE oscillators, it is useful to retain the general inequality. Note that this synchronization between node 1 and 2 occurs regardless of the dynamics of all the other oscillators, which in particular do not need to be synchronized and can have arbitrary frequencies. We stress that this is a consequence of the topological equivalence (4); without it, $h_{1}$ and $h_{2}$ do not cancel out in the manifold $\theta_{1}=\theta_{2}$ in Eq. (2). We can even imaging to modify the inner part of the remainder of the graph by dynamically removing, adding, or rewiring some of its links, as well as by allowing for the presence of any site-dependent noise: as far as such links are not those arriving at nodes 1 and 2, and as far as such noise applies only to the other nodes, the sufficient condition $k^{(\mathrm{OUT})}\leq 2(1-\gamma)$ remains satisfied, i.e., the subsystem fixed point $\theta_{1}=\theta_{2}$ keeps being stable.
@@
Equal frequencies with $a_{1,2}=0$. Before analyzing the most general case, it is worth also considering the sub-case in which the two nodes are disconnected,i.e., $a_{1,2}=0$ and have equal frequencies. In this case, Eq. (7) is of no help because $\theta_{1}=\theta_{2}$ is no longer a stable attractor, in other words, in general, the two oscillators do not get synchronized. However, from Eq. (1) we see that, if nodes 1 and 2 are TE, $\theta_{1}$ and $\theta_{2}$ obey the same identical equation: we simply have that one oscillator follows the other along the same trajectory and, in particular, their phases will remain equal at any instant in the case of same initial conditions. In this specific case, ${\Theta}=0$ is no longer an attractor but rather a constant of the motion.
@@
Different frequencies with $a_{1,2}=1$. Let us now allow for the two frequencies to be different. If the pair is isolated, i.e., $k^{(\mathrm{OUT})}=0$, we have $h^{\prime}(\theta^{*})=0$ and Eq. (7) returns the known critical condition for the synchronization of ${N=}N^{\prime}=2$ oscillators: $2J>|\omega_{2}-\omega_{1}|$, the stability condition of the fixed point $\Theta$} being $\cos({\Theta})>0$. In the general case, $h^{\prime}(\theta^{*})\neq 0$, although Eq. (7) admits a formal fixed point, the fact that $\theta^{*}$ depends on $\theta_{1}$ (besides $\Theta$}), which in turn is a function of time, makes this formal fixed point useless. Our approach here is different: we assume that both $|\omega_{2}-\omega_{1}|/J$ and the initial condition $|{\Theta}_{0}|=|\theta_{2}(t=0)-\theta_{1}(t=0)|$ are sufficiently small and seek a bounding solution by exploiting again the bounds (8) and (10). Let us suppose ${\Omega}=\omega_{2}-\omega_{1}>0$ and $0<{\Theta}_{0}<1$. By continuity, at small enough times, we also have $0<{\Theta}<1$ and from Eq. (7) we get
@@
Eq. (14) implies that, if $\lambda=k^{(\mathrm{OUT})}-2(1-\gamma)<0$, $\Theta$}, with initial condition $0<{\Theta}_{0}<1$, remains bounded as
@@
Note that, for $0<\Omega/J|\lambda|<\Theta_{0}$, all the above procedure turns out to be consistent with the required bound $0<{\Theta}<1$ for any $t$.
In conclusion, we have proven that, under the three conditions $k^{(\mathrm{OUT})}<2(1-\gamma)$ (equivalent to $k^{(\mathrm{OUT})}\leq 1$), $0<{\Theta}_{0}<1$, and $0<\Omega/J|\lambda|<\Theta_{0}$, we have $0\leq\Theta\leq\Omega/J|\lambda|$ when $t\to\infty$. In other words, the phases of the two oscillators get asymptotically close to each other and the larger is $J$ the closer they stay. Moreover, as in the case of equal frequencies, we see that this bound holds regardless of the status of all the other oscillators located on the remainder of the graph and, again, any change in it, cannot affect the synchronization of the pair.
@@
Generalization to $N^{\prime}>2$ TE oscillators. Let us now consider a subsystem of $N^{\prime}=3$ TE nodes, say nodes 1, 2 and 3. Besides being TE with respect to the remainder of the graph, we want them to be TE among each other; in particular, each of them must have the same number of links pointing to the other two TE nodes. For $N^{\prime}=3$ there exists only one possibility, the one where $k^{(\mathrm{IN})}=2$, i.e., the three nodes form a triangle. As anticipated in the Introduction, $k^{(\mathrm{IN})}$ is the (common) number of links emanating from each node of the group and pointing to the other nodes of the same group. Note that, formally, both the subsystem $N^{\prime}=3$ and the already seen case $N^{\prime}=2$ (where it was $k^{(\mathrm{IN})}=1$), are FC graphs with $N^{\prime}$ nodes. However, as Fig. 1 shows, when $N^{\prime}>3$, there exist more configurations in which the $N^{\prime}$ TE nodes can be arranged. In fact, the number of possible arrangements tends to grow exponentially with $N^{\prime}$, but with a smaller rate for $N^{\prime}$ odd (see also SM ).
Figure 1: Graph associated to model (1) with $N=22$ nodes and $L=46$ links (15 red-ticker and 31 black-thin). The graph contains five groups of TE nodes: one FC with $N^{\prime}=3$ (nodes 1, 2 and 3), two FC with $N^{\prime}=2$ (nodes 7 and 6, and nodes 15 and 16, respectively); one with $N^{\prime}=4$ forming a square (nodes 9, 10, 11, and 12); and one FC with $N^{\prime}=4$ (nodes 19, 20, 21, 22). In each group, the internal links are drawn as ticker (red). In the four FC groups we have $k^{(\mathrm{OUT})}\leq k^{(\mathrm{IN})}$ while in the other group we have $k^{(\mathrm{OUT})}<k^{(\mathrm{IN})}$, hence, in each group, the TC for synchronization are satisfied.
We introduce some notation and point out a few crucial points. Given $N^{\prime}$ TE nodes with indices $1,2,\ldots,N^{\prime}$, we indicate their corresponding phase differences by ${\Theta}_{i,j}=\theta_{i}-\theta_{j}$ and frequency differences by ${\Omega}_{i,j}=\omega_{i}-\omega_{j}$. Observe that ${\Theta}_{j,i}=-{\Theta}_{i,j}$ and that the variables $\{{\Theta}_{i,j}\}$ with $j>i$, are not all independent. For example, for $N^{\prime}=3$ we have the constrain ${\Theta}_{3,1}={\Theta}_{3,2}+{\Theta}_{{2,1}}$. In general, given $N^{\prime}$, we can always write a system of $N^{\prime}-1$ independent equations involving $N^{\prime}-1$ independent variables. We shall also assume $0<{\Theta}_{i,j}(0)<1/(N^{\prime}-1)$ so that, at small enough times, we also have $0<{\Theta}_{i,j}(t)<1/(N^{\prime}-1)$. Our general strategy is to use the bounds (8), (10); if the found bounding solution satisfies the above constrains for any $t$, the procedure is consistent. The resulting bounding system can be written vectorially:
@@
where $\bm{{\Theta}}=({\Theta}_{2,1},{\Theta}_{3,2},\ldots,{\Theta}_{N^{\prime},N^{\prime}-1})^{T}$, $\bm{{\Omega}}=({\Omega}_{2,1},{\Omega}_{3,2},\ldots,{\Omega}_{N^{\prime},N^{\prime}-1})^{T}$, and $\bm{B}$ is a $(N^{\prime}-1)\times(N^{\prime}-1)$ matrix that depends on the parameters $N^{\prime}$, $k^{(\mathrm{OUT})}$, $k^{(\mathrm{IN})}$, and $\gamma$. Let $\lambda_{i}$ and $\bm{u}_{i}$ be the eigenvalues and normalized eigenvectors of $\bm{B}$, respectively. The coefficients $c_{i}(t)=\bm{{\Theta}}\cdot\bm{u}_{i}$ satisfy the decoupled system $dc_{i}(t)/dt\leq\bm{{\Omega}}\cdot\bm{u}_{i}+J\lambda_{i}c_{i}(t)$. As a consequence, each component of the vector $\bm{{\Theta}}$ remains bounded if all the eigenvalues of $\bm{B}$ are negative, leading to the TC (note that the $\lambda$’s do not depend on $J$ or $\bm{{\Omega}}$). For example, in the FC case $N^{\prime}=3$ we have